If HJ = x - 1, and DF = 6x -18, then DF =  _______


1
Expert's answer
2021-02-24T07:25:08-0500

"HJ = x-1"

"DF = 6x-18 = 6*(x-1)-12 = 6HJ-12=6(HJ-2)"

Answer:"DF =6(HJ-2)"